OpenStudy (alystar):

Let (𝑥) = 𝑥 − 2 𝑎𝑛𝑑 𝑔(𝑥) = 𝑥^2 . Find 𝑔(𝑓(−3))

OpenStudy (alystar):

a) 25 b) 7 c) -9 d) -45

OpenStudy (peachpi):

work from the inside out. First find f(-3) by plugging in -3 for x in f(x)

OpenStudy (alystar):

I'm not entirely sure where to plug it in.

OpenStudy (peachpi):

plug it in for x f(x) = x - 2 f(-3) = -3 - 2

OpenStudy (alystar):

so -5

OpenStudy (peachpi):

yes. now plug in -5 for x in g(x) g(x) = x^2 g(-5) = (-5)^2

OpenStudy (alystar):

25, so the answer is A

OpenStudy (peachpi):

right

OpenStudy (alystar):

Thanks, "𝑔(𝑓(−3))" just threw me off.